OpenStudy (nuccioreggie):

A circle is cut from a square piece of cloth, as shown: A square, one side labeled as 36 inches, has a circle inside it. The circle touches all the sides of the square. The portion of the square outside the circle is shaded. How many square inches of cloth are cut from the square? (π = 3.14) 1,017.36 in2 1,182.96 in2 1,276.00 in2 1,489.24 in2

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

I think its B

OpenStudy (michele_laino):



OpenStudy (nuccioreggie):

It a?

OpenStudy (nuccioreggie):

What is the correct answer choice?

OpenStudy (michele_laino):

the area of shaded region, is: 36-(3.14*18*18)=...

OpenStudy (michele_laino):

oops.. area= 36^2- (3.14*18*18)=1296-(3.14*18*18)=...

OpenStudy (nuccioreggie):

278.64

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

now your answer is the area of the circle, which is: 3.14*18*18=...

OpenStudy (nuccioreggie):

A

OpenStudy (michele_laino):

that's right!